OpenStudy (ashchu117):

I WILL GIVE MEDALS!!!!!!!! Solve for x: 2 over 5 (x - 4) = 2x. A. -8 B. negative 1 over 2 C. 1 D. -1

sammixboo (sammixboo):

Well you can plug in your answer choices as x, and then solve your equation. If both sides are not equal, then that is not your answer

sammixboo (sammixboo):

One way of doing it.

jimthompson5910 (jim_thompson5910):

I'll get you started \[\Large \frac{2}{5}(x - 4) = 2x\] \[\Large 5*\frac{2}{5}(x - 4) = 5*2x\] \[\Large 2(x - 4) = 10x\] \[\Large 2x - 8 = 10x\] and let you take over from here

OpenStudy (ashchu117):

8x = 8?

jimthompson5910 (jim_thompson5910):

close but no

sammixboo (sammixboo):

Nice explaining! I am not really in the thinking mood ; _ ; nor helping mood D: I hate it when I am like that

OpenStudy (ashchu117):

so its C???

jimthompson5910 (jim_thompson5910):

that's fine sammixboo it should be 2x - 8 = 10x -8 = 10x - 2x -8 = 8x x = ???

OpenStudy (ashchu117):

x = 1- ?

OpenStudy (ashchu117):

D?

jimthompson5910 (jim_thompson5910):

you mean x = -1 right?

OpenStudy (ashchu117):

lol yeah XD im so tirededefian9fjashfeqfj9q8uhr3

jimthompson5910 (jim_thompson5910):

yep correct

jimthompson5910 (jim_thompson5910):

Check: \[\Large \frac{2}{5}(x - 4) = 2x\] \[\Large \frac{2}{5}(-1 - 4) = 2(-1)\] \[\Large \frac{2}{5}(-5) = 2(-1)\] \[\Large -2 = -2\] so that verifies the solution

OpenStudy (ashchu117):

thx bro/broett(female bro)!!!!!!

jimthompson5910 (jim_thompson5910):

np